STATISTICS AND PROBABILITY QUESTION:

The computers of nine engineers at a certain company are to be replaced. Four of the engineers have selected laptops and the other 5 have selected desktops. Suppose that four computers are randomly selected (a) How many different ways are there to select four of the nine computers to be set up? A. 126 B. 70 C. 40320 (b) What is the probability that exactly three of the selected computers are desktops? A. 0.2289 B. 0.4286 C. 0.3175 (c) What is the probability that at least three desktops are selected? A. 0.1667 B. 0.5000 C. 0.1190 D. 210 E. 56 D. 0.1143 E. 0.158 D. 0.2477 E. 0.3571
Answer #1

a)

number of ways =9C4 =126

b)

P(three are desktops)=5C3*4C1/9C4 = 0.3175

option C

c)P(at least 3) =P(X=3)+P(X=4)=5C3*4C1/9C4 +5C44C0/9C4 =0.3571

option E